# Testing the hypothesis of vector X17 boson by $D$ meson, Charmonium, and $\phi$ meson decays

### v2

@@ II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
We consider the $1^{-}\rightarrow 0^{-}$ heavy meson transitions $H^{*}\rightarrow He^{+}e^{-}$. The Lorentz-invariant matrix elements, accounting for the contribution by the intermediate photon and X17 boson, are denoted by $-i\mathcal{M}^{\gamma}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ and $-i\mathcal{M}^{X}\left(H^{*}\rightarrow He^{+}e^{-}\right)$, respectively. We have
$-i\mathcal{M}^{\gamma}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ $=$ $T^{\gamma}_{\mu}\frac{-ig^{\mu\nu}}{q^{2}+i\epsilon}(-ie)\bar{u}\left(p_{-}\right)\gamma_{v}v\left(p_{+}\right),$
$-i\mathcal{M}^{X}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ $=$ $T^{X}_{\mu}\frac{i\left(-g^{\mu\nu}+q^{\mu}q^{v}/m_{X}^{2}\right)}{q^{2}-m_{X}^{2}+im_{X}\Gamma_{X}}(-ie\varepsilon_{e})\bar{u}\left(p_{-}\right)\gamma_{v}v\left(p_{+}\right),$ (2)
@@
Figure 1: The experimentally measured as well as theoretically predicted values for $R_{D_{s}^{*+}}$. Various theoretical predictions are distinguished by different values taken for $\varepsilon_{c}$ and $\varepsilon_{s}$. VMDa stands for the calculation of Castro:2021gdf using VMD approach and employing $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ extracted by Feng:2016jff ; Fornal:2017msy . VMDa Corrected Result refers to our recalculation of $R_{D_{s}^{*+}}$ using the same values of $\varepsilon_{c}$ and $\varepsilon_{s}$. VMDb denoted in blue represents our result of using $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays, and that denoted in black is obtained by employing $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The SM prediction is denoted as SM__\__ Photon.
@@
where $\epsilon_{\mu}$ is the polarization of the vector meson. Using the current data of $V\to e^{+}e^{-}$ ParticleDataGroup:2024cfk , it is found that $(f_{\rho^{0}},f_{\omega},f_{\phi})=(0.171,0.155,0.232)\ {\rm GeV}^{2}$ Castro:2021gdf . Finally the parameter $\lambda$ can be determined by the data of semi-leptonic decays $D\to\pi(K)l\nu_{l}$ and $D\to K^{*}l\nu_{l}$ as pointed out in Colangelo:1993zq . With the updated data, it is found that Castro:2021gdf
@@
We point out that our $m_{s}(q^{2})^{-1}$ agrees with that of Castro:2021gdf . On the other hand, $m_{u,d}(q^{2})^{-1}$ given by Eq. (5) of Castro:2021gdf is twice larger than ours. It is easily seen that our results in Eq. (10) satisfy SU(3) symmetry relation $m_{u}(q^{2})^{-1}=m_{d}(q^{2})^{-1}=m_{s}(q^{2})^{-1}$ in the SU(3) symmetry limit $f_{\rho^{0}}=f_{\omega}=f_{\phi}$, and $m_{\rho^{0}}=m_{\omega}=m_{\phi}$.
@@
Fig. 1 presents the experimentally measured value of $R_{D^{*+}_{s}}$ as well as the corresponding theoretical predictions with and without contributions from the X17 boson. The measured value of $R_{D^{*+}_{s}}$ is marked in red while the SM prediction of this ratio is marked in yellow, which considers only the photon contribution and gives $R_{D^{*+}_{s}}^{\gamma}=6.7\times 10^{-3}$. This is consistent with the experimental data, $R_{D^{*+}_{s}}^{\text{exp}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . Including the contribution by X17 boson with $\varepsilon_{u}=$$\pm 3.7\times 10^{-3},\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff , we obtain $R_{D^{*+}_{s}}=(2.6\pm 1.4)\times 10^{-2}$ as denoted by “VMDa Corrected Result”, which is in a mild tension with the data. For current and subsequent calculations, we take $m_{X}=17$ MeV. It is found that the decay ratios, such as $R_{D^{*+}_{s}}$ and those involving other mesons, are not sensitive to $m_{X}$ over its allowed range. Our result disagrees with the calculation by Castro:2021gdf which gives $R_{D_{s}^{*+}}=(9.82\pm 0.60)\times 10^{-3}$. The detailed comparison between two results will be given in the next paragraph. The results denoted by VMDb are obtained with $\varepsilon_{u,d}$ taken from Denton:2023gat . They are divided into two cases. The first case, denoted in blue, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for ${}^{8}{\rm Be}^{*}$ nuclear decays. The second case, denoted in black, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered. The uncertainties for theoretical predictions arise from the uncertainty of coupling parameter $\lambda$. It is seen that both predictions under VMDb are consistent with the experimental data.
@@
To compare with Ref. Castro:2021gdf in details, we note that our calculations reproduce $\varepsilon_{Q}/m_{H^{*}}=1.75\times 10^{-3}\ {\rm GeV}^{-1}$ and $\varepsilon_{q}/m_{q}(m_{X}^{2})=-7.83\times 10^{-3}\ {\rm GeV}^{-1}$ for the decay $D_{s}^{*+}\to D_{s}^{+}X$ given in Table II of that paper. Hence $F_{H^{*}HX}(m_{X}^{2})$ defined by Eq. (9), is easily obtained as $(-0.63\pm 0.05)\times 10^{-2}\ {\rm GeV}^{-1}$ rather than $(-0.91\pm 0.06)\times 10^{-2}\ {\rm GeV}^{-1}$ given in Table II of Castro:2021gdf . Here the uncertainty of $F_{H^{*}HX}(m_{X}^{2})$ is due to the uncertainty of $\lambda$. Furthermore, the quantity $e_{q}/m_{q}(0)$ in $D_{s}^{*+}\to D_{s}^{+}\gamma$ is related to $\varepsilon_{q}/m_{q}(m_{X}^{2})$ by a simple scaling: $e_{q}/m_{q}(0)=e_{q}/\varepsilon_{q}\cdot\varepsilon_{q}/m_{q}(0)\simeq e_{q}/\varepsilon_{q}\cdot\varepsilon_{q}/m_{q}(m_{X}^{2})$. With $e_{s}=-1/3$, $\varepsilon_{s}=-7.4\times 10^{-3}$, and $\varepsilon_{s}/m_{s}(m_{X}^{2})=-7.83\times 10^{-3}\ {\rm GeV}^{-1}$ as just mentioned, one obtains $e_{s}/m_{s}(0)=-0.35\ {\rm GeV}^{-1}$ rather than $-0.48\ {\rm GeV}^{-1}$ given in Table I of  Castro:2021gdf . Since $e_{c}/m_{D^{*+}_{s}}=0.32\ {\rm GeV}^{-1}$, one obtains $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)=-0.038\pm 0.020\ {\rm GeV}^{-1}$ where the uncertainty again arises from the uncertainty of $\lambda$. We observe that the central value of $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)$ is much smaller than that in Castro:2021gdf which gives $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)=-0.17\pm 0.03\ {\rm GeV}^{-1}$. Hence we predict a much larger $R^{X}_{D^{*+}_{s}}$ due to a much smaller $F^{2}_{D^{*+}_{s}D^{+}_{s}\gamma}(0)$ on the right hand side of Eq. (7). This leads to a much larger overall ratio $R_{D^{*+}_{s}}$.
@@
$D^{\ast 0}\rightarrow D^{0}e^{+}e^{-}$ $6.7\times 10^{-3}$ $5.6\times 10^{-7}$ $6.7\times 10^{-3}$ $(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq
$D^{\ast+}\rightarrow D^{+}e^{+}e^{-}$ $6.6\times 10^{-3}$ $(1.2\pm 0.4)\times 10^{-3}$ $(7.8\pm 0.4)\times 10^{-3}$
$D_{s}^{\ast+}\rightarrow D_{s}^{+}e^{+}e^{-}$ $6.7\times 10^{-3}$ $(4.2\pm 3.1)\times 10^{-3}$ $(1.1\pm 0.3)\times 10^{-2}$ $(7.2_{-1.6}^{+1.8})\times 10^{-3}$ CLEO:2011mla
We consider a similar analysis for the decay $D^{*0}\rightarrow D^{0}e^{+}e^{-}$. The experimental measurement gives $R^{\text{exp}}_{D^{*0}}=(11.08\pm 0.9)\times 10^{-3}$ BESIII:2021vyq and it is marked in red in Fig. 2. The SM contribution due to the photon mediation yields $R_{D^{*0}}^{\gamma}=6.7\times 10^{-3}$. The full contributions $R_{D^{*0}}\equiv R_{D^{*0}}^{\gamma}+R_{D^{*0}}^{X}$ corresponding to different ranges of new physics couplings are labeled as VMDa and VMDb, respectively. Table 1 presents results of $R_{H^{*}}$ arising from intermediate photon and X17 boson contributions for three $D^{*}$ meson decay modes within the VMD model, employing $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in ${}^{8}{\rm Be}^{*}$ decays considered Denton:2023gat . Considering isospin mixing and breaking effects in ${}^{8}{\rm Be}^{*}$ decays, i.e., $\varepsilon_{u}=\varepsilon_{c}=\pm 9.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.5\times 10^{-3}$, the corresponding $R_{H^{*}}$ values are $6.7\times 10^{-3}$, $(7.4\pm 0.3)\times 10^{-3}$, and $(9.3\pm 1.9)\times 10^{-3}$ for $D^{*0}$, $D^{*+}$, and $D^{*+}_{s}$ decays, respectively.
@@
Figure 2: The experimentally measured and theoretically predicted values for $R_{D_{0}^{*}}$. VMDa is the result obtained by VMD approach where $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff ; Fornal:2017msy . The result of VMDb denoted in blue is calculated by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$, and that of VMDb denoted in black is obtained by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. The SM prediction is denoted as SM__\__ Photon.
We observe that the above adjustments of $\varepsilon_{c}$ and $\varepsilon_{s}$ naturally affect the decays of Charmonium ($c\bar{c}$) and $\phi$ ($s\bar{s}$) mesons. Hence the determination of these coupling parameters should involve the data of these decay channels.

### v3

@@ II X17 contributions to heavy meson decays $H^{*}\to H+e^{+}+e^{-}$
We consider the $1^{-}\rightarrow 0^{-}$ heavy meson transitions $H^{*}\rightarrow He^{+}e^{-}$. The Lorentz-invariant matrix elements, accounting for the contribution by the intermediate photon and X17 boson, are denoted by $i\mathcal{M}^{\gamma}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ and $i\mathcal{M}^{X}\left(H^{*}\rightarrow He^{+}e^{-}\right)$, respectively. We have
$-i\mathcal{M}^{\gamma}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ $=$ $T^{\gamma}_{\mu}\frac{-ig^{\mu\nu}}{q^{2}+i\epsilon}(-ie)\bar{u}\left(p_{-}\right)\gamma_{\nu}v\left(p_{+}\right),$
$-i\mathcal{M}^{X}\left(H^{*}\rightarrow He^{+}e^{-}\right)$ $=$ $T^{X}_{\mu}\frac{i\left(-g^{\mu\nu}+q^{\mu}q^{\nu}/m_{X}^{2}\right)}{q^{2}-m_{X}^{2}+im_{X}\Gamma_{X}}(-ie\varepsilon_{e})\bar{u}\left(p_{-}\right)\gamma_{\nu}v\left(p_{+}\right),$ (2)
@@
Figure 1: The experimentally measured as well as theoretically predicted values for $R_{D_{s}^{*+}}$. Various theoretical predictions are distinguished by different values taken for $\varepsilon_{c}$ and $\varepsilon_{s}$. VMDa stands for the calculation of Castro:2021gdf using VMD approach and employing $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ extracted by Feng:2016jff ; Fornal:2017msy . VMDa Corrected Result refers to our recalculation of $R_{D_{s}^{*+}}$ using the same values of $\varepsilon_{c}$ and $\varepsilon_{s}$. VMDb denoted in blue represents our result of using $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for $^{8}{\rm Be}^{*}$ nuclear decays, and that denoted in black is obtained by employing $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered Denton:2023gat . The SM prediction is denoted as SM_\_ Photon.
@@
where $\epsilon_{\mu}$ is the polarization of the vector meson. Using the current data of $V\to e^{+}e^{-}$ ParticleDataGroup:2024cfk , it is found that $(f_{\rho^{0}},f_{\omega},f_{\phi})=(0.171,0.155,0.232)\ {\rm GeV}^{2}$ Castro:2021gdf . Finally, the parameter $\lambda$ can be determined by the data of semi-leptonic decays $D\to\pi(K)l\nu_{l}$ and $D\to K^{*}l\nu_{l}$ as pointed out in Colangelo:1993zq . With the updated data, it is found that Castro:2021gdf
@@
We point out that our $m_{s}(q^{2})^{-1}$ agrees with that of Castro:2021gdf . On the other hand, $m_{u,d}(q^{2})^{-1}$ given by Eq. (5) of Castro:2021gdf is twice larger than ours. To see how this difference propagates into the final prediction, let us define the heavy and light contributions as $H_{\gamma}\equiv e_{Q}/m_{H^{*}}$, $L^{(q)}_{\gamma}(q^{2})\equiv e_{q}/m_{q}(q^{2})$, $H_{X}\equiv\varepsilon_{Q}/m_{H^{*}}$, and $L^{(q)}_{X}(q^{2})\equiv\varepsilon_{q}/m_{q}(q^{2})$. Then, Eqs. (7) and (9) imply
$\frac{F_{H^{*}HX}(q^{2})}{F_{H^{*}H\gamma}(0)}=\frac{H_{X}+L^{(u,d)}_{X}(q^{2})}{H_{\gamma}+L^{(u,d)}_{\gamma}(0)}\,\quad\Rightarrow\quad R^{X}_{H^{*}}\propto\left|\frac{H_{X}+L_{X}}{H_{\gamma}+L_{\gamma}}\right|^{2}.$ (14)
The factor of 2 in Ref. Castro:2021gdf enters through the light-quark terms in both the numerator and the denominator of the ratio $R^{X}_{H^{*}}$ or specifically $R^{X}_{D^{*}}$. This corresponds to the replacement $L^{(u,d)}_{\gamma,X}\to 2L^{(u,d)}_{\gamma,X}$ in Eq. (14). The total ratio $R_{D^{*}}\equiv R^{\gamma}_{D^{*}}+R^{X}_{D^{*}}$ is therefore affected. The exact SU(3) symmetry limit amounts to setting $f_{\rho}=f_{\omega}=f_{\phi}$ and $m_{\rho}=m_{\omega}=m_{\phi}$. This naturally leads to another SU(3) symmetry relation $m_{u}^{-1}(q^{2})=m_{d}^{-1}(q^{2})=m_{s}^{-1}(q^{2})$ by applying the first two relations to Eq. (10).
Fig. 1 presents the experimentally measured value of $R_{D^{*+}_{s}}$ as well as the corresponding theoretical predictions with and without contributions from the X17 boson. The measured value of $R_{D^{*+}_{s}}$ is marked in red while the SM prediction of this ratio is marked in yellow, which considers only the photon contribution and gives $R_{D^{*+}_{s}}^{\gamma}=6.5\times 10^{-3}$. This is consistent with the experimental data, $R_{D^{*+}_{s}}^{\text{exp}}=\left(7.2_{-1.6}^{+1.8}\right)\times 10^{-3}$ CLEO:2011mla . We note that such a consistency does not render this channel irrelevant to the X17 hypothesis. The X17 boson is motivated by anomalous nuclear transitions of excited $^{8}{\rm Be}$, $^{4}{\rm He}$ and $^{12}{\rm C}$. Clearly, any viable X17 scenario must also be consistent with the data of electromagnetic Dalitz-type transitions in heavy-meson systems. The absence of a deviation to the SM predicted $R_{D_{s}^{*+}}$ in CLEO data CLEO:2011mla  [^2] The CLEO analysis CLEO:2011mla focuses on the integrated rate $R_{D_{s}^{*+}}$ and the distribution of $e^{+}e^{-}$ invariant mass $M_{ee}$. Since the latter distribution still contains large uncertainties, we only use the decay ratio $R_{D_{s}^{*+}}$ to test the X17 hypothesis in $D_{s}^{*+}\to D^{+}_{s}e^{+}e^{-}$ channel. therefore imposes a strong constraint on any new physics contribution to $H^{*}\to He^{+}e^{-}$. This is precisely our main motivation for performing this cross-channel test for X17 hypothesis.
@@
By including the contribution from X17 boson with $\varepsilon_{u}=$ $\pm 3.7\times 10^{-3},\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff , we obtain $R_{D^{*+}_{s}}=(2.6\pm 1.4)\times 10^{-2}$ as denoted by “VMDa Corrected Result”, which is in a mild tension with the data. For current and subsequent calculations, we take $m_{X}=16.85$ MeV. It is found that the decay ratios, such as $R_{D^{*+}_{s}}$ and those involving other mesons, are not sensitive to $m_{X}$ over its allowed range. Our result disagrees with the calculation by Castro:2021gdf which gives $R_{D_{s}^{*+}}=(9.82\pm 0.60)\times 10^{-3}$. The detailed comparison between two results will be given in the next paragraph. The results denoted by VMDb are obtained with $\varepsilon_{u,d}$ taken from Denton:2023gat . They are divided into two cases. The first case, denoted in blue, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$ which does not consider isospin effects for $^{8}{\rm Be}^{*}$ nuclear decays. The second case, denoted in black, is obtained by taking $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$ where both isospin mixing and breaking effects are considered. The uncertainties for theoretical predictions arise from the uncertainty of coupling parameter $\lambda$. It is seen that both predictions under VMDb are consistent with the experimental data.
@@
To compare with Ref. Castro:2021gdf in details, we note that our calculations reproduce $\varepsilon_{Q}/m_{H^{*}}=1.75\times 10^{-3}\ {\rm GeV}^{-1}$ and $\varepsilon_{q}/m_{q}(m_{X}^{2})=-7.83\times 10^{-3}\ {\rm GeV}^{-1}$ for the decay $D_{s}^{*+}\to D_{s}^{+}X$ given in Table II of that paper. Hence, $F_{H^{*}HX}(m_{X}^{2})$ defined by Eq. (9), is easily obtained as $(-0.63\pm 0.05)\times 10^{-2}\ {\rm GeV}^{-1}$ rather than $(-0.91\pm 0.06)\times 10^{-2}\ {\rm GeV}^{-1}$ given in Table II of Castro:2021gdf . Here, the uncertainty of $F_{H^{*}HX}(m_{X}^{2})$ is due to the uncertainty of $\lambda$. Furthermore, the quantity $e_{q}/m_{q}(0)$ in $D_{s}^{*+}\to D_{s}^{+}\gamma$ is related to $\varepsilon_{q}/m_{q}(m_{X}^{2})$ by a simple scaling: $e_{q}/m_{q}(0)=e_{q}/\varepsilon_{q}\cdot\varepsilon_{q}/m_{q}(0)\simeq e_{q}/\varepsilon_{q}\cdot\varepsilon_{q}/m_{q}(m_{X}^{2})$. With $e_{s}=-1/3$, $\varepsilon_{s}=-7.4\times 10^{-3}$, and $\varepsilon_{s}/m_{s}(m_{X}^{2})=-7.83\times 10^{-3}\ {\rm GeV}^{-1}$ as just mentioned, one obtains $e_{s}/m_{s}(0)=-0.35\ {\rm GeV}^{-1}$ rather than $0.48\ {\rm GeV}^{-1}$ given in Table I of  Castro:2021gdf . Since $e_{c}/m_{D^{*+}_{s}}=0.32\ {\rm GeV}^{-1}$, one obtains $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)=-0.038\pm 0.020\ {\rm GeV}^{-1}$ where the uncertainty again arises from the uncertainty of $\lambda$. We observe that the central value of $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)$ is much smaller than that in Castro:2021gdf which gives $F_{D^{*+}_{s}D^{+}_{s}\gamma}(0)=-0.17\pm 0.03\ {\rm GeV}^{-1}$. Hence, we predict a much larger $R^{X}_{D^{*+}_{s}}$ due to a much smaller $F^{2}_{D^{*+}_{s}D^{+}_{s}\gamma}(0)$ on the right hand side of Eq. (7). This leads to a much larger overall ratio $R_{D^{*+}_{s}}$.
@@
$D^{\ast 0}\rightarrow D^{0}e^{+}e^{-}$ $6.4\times 10^{-3}$ $5.6\times 10^{-7}$ $6.4\times 10^{-3}$ $(11.08\pm 0.90)\times 10^{-3}$ BESIII:2021vyq
$D^{\ast+}\rightarrow D^{+}e^{+}e^{-}$ $6.4\times 10^{-3}$ $(1.2\pm 0.4)\times 10^{-3}$ $(7.6\pm 0.4)\times 10^{-3}$
$D_{s}^{\ast+}\rightarrow D_{s}^{+}e^{+}e^{-}$ $6.5\times 10^{-3}$ $(4.2\pm 3.1)\times 10^{-3}$ $(1.1\pm 0.3)\times 10^{-2}$ $(7.2_{-1.6}^{+1.8})\times 10^{-3}$ CLEO:2011mla
We consider a similar analysis for the decay $D^{*0}\rightarrow D^{0}e^{+}e^{-}$. The experimental measurement gives $R^{\text{exp}}_{D^{*0}}=(11.08\pm 0.90)\times 10^{-3}$ BESIII:2021vyq and it is marked in red in Fig. 2. The SM contribution due to the photon mediation yields $R_{D^{*0}}^{\gamma}=6.4\times 10^{-3}$. The full contributions $R_{D^{*0}}\equiv R_{D^{*0}}^{\gamma}+R_{D^{*0}}^{X}$ corresponding to different ranges of new physics couplings are labeled as VMDa and VMDb, respectively. Table 1 presents results of $R_{H^{*}}$ arising from intermediate photon and X17 boson contributions for three $D^{*}$ meson decay modes within the VMD model, employing $\varepsilon_{u}=\varepsilon_{c}=\pm 5.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.9\times 10^{-3}$ that corresponds to best-fit parameters without isospin effects in $^{8}{\rm Be}^{*}$ decays considered Denton:2023gat . Considering isospin mixing and breaking effects in $^{8}{\rm Be}^{*}$ decays, i.e., $\varepsilon_{u}=\varepsilon_{c}=\pm 9.0\times 10^{-4}$ and $\varepsilon_{d}=\varepsilon_{s}=\mp 2.5\times 10^{-3}$, the corresponding $R_{H^{*}}$ values are $6.4\times 10^{-3}$, $(7.2\pm 0.3)\times 10^{-3}$, and $(9.1\pm 1.9)\times 10^{-3}$ for $^{*0}, D+D^{*+}$, and $D^{*+}_{s}$ decays, respectively.
It is seen that the SM predictions to $R^{\gamma}_{H^{*}}$ for $H^{*}=D^{*0}$, $D^{*+}$, and $D_{s}^{*+}$ as shown in Table 1 are almost identical. We stress that they do not arise from imposing SU(3) symmetry. In fact, we use physically measured vector meson masses and decay constants in our calculations instead of imposing SU(3) symmetry. To see this, we refer to Eq. (6). The prefactor $\alpha_{\rm EM}/(3\pi)$ and the $e^{+}e^{-}$ conversion kernel $(1/q^{2})\,(1+2m_{e}^{2}/q^{2})\,\sqrt{1-4m_{e}^{2}/q^{2}}$ are identical for all channels. Accordingly, any channel-to-channel differences can arise only from (i) the value of the integration upper limit $q_{\max}^{2}\equiv(m_{H^{*}}-m_{H})^{2}$, (ii) the form-factor shape $F_{H^{*}H\gamma}^{2}(q^{2})/F_{H^{*}H\gamma}^{2}(0)$, and (iii) the kinematic factor $K(q^{2})\equiv\lambda^{3/2}\!\left(m_{H^{*}}^{2},m_{H}^{2},q^{2}\right)/\left(m_{H^{*}}^{2}-m_{H}^{2}\right)^{3}$ with $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2xy-2xz-2yz$. First, the value for $q_{\max}^{2}$ only differs by a few percent among $D^{*+}$, $D^{*0}$, and $D_{s}^{*+}$ ParticleDataGroup:2024cfk . Second, with $q_{\max}^{2}\sim 0.02\ {\rm GeV}^{2}$ and $m_{V}^{2}\sim 0.6$$1.0~\mathrm{GeV}^{2}$ ParticleDataGroup:2024cfk , $q^{2}/m_{V}^{2}\leq q_{\max}^{2}/m_{V}^{2}\ll 1$. This implies $F_{H^{*}H\gamma}(q^{2})/F_{H^{*}H\gamma}(0)=1+\mathcal{O}\!\left(\ q^{2}/m_{V}^{2}\right)$, i.e., the SU(3) flavor-dependent contributions are of the order $q^{2}/m_{V}^{2}$, which is in the range of $(2-3)\%$. Hence, the SU(3) breaking effects here are also at the same level. Third, the kinematic factor $K(q^{2})$ is a monotonic decreasing function of $q^{2}$ and its value at $q^{2}=q^{2}_{\rm max}$ vanishes. $K(q^{2})$ therefore remains positive definite in the allowed kinematic range. Taking $q^{2}=0.01$ GeV2 as an illustrative example, we find $K(q^{2})=0.358,\ 0.349$, and 0.329 for $H^{*}=D^{*0}$, $D^{*+}$, and $D_{s}^{*+}$, respectively. The differences among these factors are no more than $8\%$. For other $q^{2}$ values, the SU(3) breaking effects remain small. In summary, we have elaborated why SU(3) breaking effects are small in $R^{\gamma}_{H^{*}}$ for $H^{*}=D^{*0}$, $D^{*+}$, and $D_{s}^{*+}$.
@@
Figure 2: The experimentally measured and theoretically predicted values for $R_{D_{0}^{*}}$. VMDa is the result obtained by VMD approach where $\varepsilon_{c}=\varepsilon_{u}$ and $\varepsilon_{s}=\varepsilon_{d}$ with $\varepsilon_{u}=\pm 3.7\times 10^{-3}$ and $\varepsilon_{d}=\mp 7.4\times 10^{-3}$ Feng:2016jff ; Fornal:2017msy . The result of VMDb denoted in blue is calculated by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 5\times 10^{-4},\mp 2.9\times 10^{-3})$, and that of VMDb denoted in black is obtained by applying $(\varepsilon_{u},\varepsilon_{d})=(\pm 9\times 10^{-4},\mp 2.5\times 10^{-3})$. The SM prediction is denoted as SM_\_ Photon.
We observe that the above adjustments of $\varepsilon_{c}$ and $\varepsilon_{s}$ naturally affect the decays of Charmonium ($c\bar{c}$) and $\phi$ ($s\bar{s}$) mesons. Hence, the determination of these coupling parameters should involve the data of these decay channels.